Calculate each of the unknown currents for this circuit

uploaded imageCalculate each of the unknown currents I1, I2, and I3 for thecircuit.

I1 =
I2 =
I3 =
Answer #1

For this solution, Kirchhoff's Rules will be applied. There are three unknowns so three equations must be set up and then do a lot ofsubstitution.

Equation 1 will be the junction equation.

I1 = I2 + I3 (equation 1)

Equation 2 is chosen to loop clockwise in the upper loop (Remember that in any loop, the total potential difference must be zero). Refer toKirchhoff's rule if you do not understand why I used + and - where I did.

24 - I1R1 - 4I1 - 3I3 = 0 or 24 - 2.4I1 - 4 I1 - 3I3 = 0

Simplify to get 24 - 6.4I1 - 3I3 = 0 (equation 2)

Equation 3 is chosen to loop clockwise in the lower loop. (Same rules apply)

V +3I3 - R2I2 - 5I2 = 0 or 10 + 3I3 - 1.3I2 - 5I2 = 0

Simplify toget10 + 3I3 - 6.3I2 = 0 (equation 3)

Now to start substituting...

We know that I1 = I2 + I3 from equation 1. I will substitute that into equation 2

24 - 6.4(I2 + I3) - 3I3 = 0

Distribute and simplify to get

24- 6.4I2 - 9.4I3 = 0 (call this equation 4)

From equation 3, I will solve for I2

6.3I2 = 10 + 3I3 or, after dividing both sides by 6.3, you getI2 = 1.59 +.48I3 (Call this equation 5)

I will take equation 5 and substitute what I know for I2 into equation 4 and therefore be left with only one unknown (I3)

24 - 6.4(1.59 + .48I3) - 9.4I3 = 0

Distribute and simplify to get

13.82 -12.48I3 = 0

Solve for I3, I3 = 1.11 A

Substitute I3 back into equation 4 to find I2

24 - 6.4I2 - (9.4)(1.11) = 0

Solve for I2, I2 = 2.12 A

Finally substitute I2 and I3 back into equation 1 to find I1

I1 = 3.23 A

answered by: AMAR JYOTI